OpenStudy (imamermaid29):

Which equation represents the graph of the linear function? A. y = 4x + 4 B. y = 4x − 4 C. y = −4x + 4 D. y = −4x − 4

OpenStudy (imamermaid29):

@YoloShroom

OpenStudy (yoloshroom):

finally, can access the probnlem lol. lemme take a look

OpenStudy (yoloshroom):

so for this one, i believe we can just plug and chug with a table of values. We look at the graph to see that, when y = -4, x = 0 and when X= 1 , y =0. So which equation satisfies this table of value? @ImaMermaid29

OpenStudy (imamermaid29):

d? @YoloShroom

OpenStudy (imamermaid29):

no nvm

OpenStudy (imamermaid29):

would it be c?

OpenStudy (yoloshroom):

It would be B, uno momento porfavor, i found a flaw in my Explanation because im trash.

OpenStudy (imamermaid29):

okay :)

OpenStudy (yoloshroom):

oh shoot, sorry iw as confsuing myself. It's B because when you plug in the values Y has to equal -4. So y=4(0)-4 or y=-4(0)+4?

OpenStudy (yoloshroom):

when you get a graph, always write up a table of values, and figure it out that way :)

OpenStudy (imamermaid29):

okay thanks so much

OpenStudy (yoloshroom):

just link me if u need me.

OpenStudy (imamermaid29):

ok :)